OpenStudy (anonymous):

how do i solve √5-9 / √5-3 ?

OpenStudy (anonymous):

\[\frac{ \sqrt{5}-9 }{ \sqrt{5}-3 }\] Rationalise the denominator.

OpenStudy (anonymous):

what do you mean @surjithayer

OpenStudy (anonymous):

\[\frac{ \sqrt{5}-9 }{\sqrt{5}-3 }\times \frac{ \sqrt{5}+3 }{\sqrt{5}+3 }=\frac{ 5+3\sqrt{5}-9\sqrt{5}-27 }{5-3^2 }=\frac{ -22-6\sqrt{5} }{ 5-9 } =?\]

OpenStudy (anonymous):

and at the end how do you solve to get the answer?

OpenStudy (anonymous):

\[=\frac{ -2\left( 11+3\sqrt{5} \right) }{-4 }=\frac{ 11 }{2 }+\frac{ 3\sqrt{5} }{ 2 }\]

OpenStudy (anonymous):

okay, hmm i think i get it..